# High-Throughput Quantification of Altermagnetic Band Splitting

### v1

@@ Abstract
Altermagnetism represents a recently established class of collinear magnetism that combines zero net magnetization with momentum-dependent spin polarization, enabled by symmetry constraints rather than spin-orbit coupling. This distinctive behavior gives rise to sizable spin splitting even in materials composed of light, earth-abundant elements, offering promising prospects for next-generation spintronics applications. Despite growing theoretical and experimental interest, the discovery of altermagnetic materials remains limited due to the complexity of magnetic symmetry and the inefficiency of conventional approaches. Here, we present a comprehensive high-throughput screening of the entire MAGNDATA database, integrating symmetry analysis with spin-polarized density functional theory (DFT) calculations to identify and characterize altermagnetic candidates. Our workflow uncovers 173 materials exhibiting significant spin splitting ($\geq 50$ meV within $\pm 3$ eV of the Fermi level), spanning both metallic and semiconducting systems. Crucially, our momentum-resolved analysis reveals that the spin splitting varies strongly across the Brillouin zone, and that the maximal splitting tends to occur away from the high-symmetry paths, a result that directly informs and guides future photoemission experiments. By expanding the catalog of known altermagnets and elucidating the symmetry-protected origins of spin splitting, this work lays a robust foundation for future experimental and theoretical advances in spintronics and quantum materials discovery.
@@ I Introduction
However, discovering new altermagnets with significant spin splitting through trial-and-error experimental approaches is inherently challenging and resource-intensive. To overcome this limitation, high-throughput computational screening, leveraging reliable first-principles density functional theory (DFT) calculations, has emerged as an efficient and powerful methodology. This approach enables the rapid and systematic exploration of large material databases for candidates exhibiting desirable characteristics. Some recent efforts have used the MAGNDATA database for this purpose (see Sec. IV for a comparison), yet these studies have largely focused on specific subclasses, such as metallic altermagnets or magnonic excitations in collinear systems, thus leaving much of the altermagnetic landscape unexplored [25, 26]. In this present work, we undertake an exhaustive and systematic high-throughput screening of the entire MAGNDATA [27] database, without constraints to specific classes or properties, aiming to identify and comprehensively characterize all potential altermagnetic candidates. Through this extensive computational investigation, we successfully discover numerous previously unreported altermagnetic materials, validate several known candidates, and provide robust theoretical frameworks for future experimental realizations. This comprehensive approach not only significantly enhances the current understanding of altermagnetic phenomena but also expands the available pool of materials for next-generation spintronic devices, highlighting the critical importance and novelty of this study.
@@ II Screening workflow
We present a rigorous two-stage high-throughput computational screening workflow for the systematic identification of candidate altermagnetic materials, summarized schematically in Fig.1. In the first stage of screening (Fig.1a), we initially retrieved over 2344 experimentally characterized magnetic materials from the MAGNDATA database [27] hosted by the Bilbao Crystallographic Server. MAGNDATA is a comprehensive and publicly available repository specializing in magnetic materials, providing crystal structures and experimentally determined magnetic configurations.
@@
These magnetic materials were subsequently analyzed using the recently developed computational tool amcheck [28]. Amcheck operationalizes the symmetry-based theoretical framework underpinning altermagnetism by assessing symmetry operations relating spin-up and spin-down magnetic sublattices. Specifically, the algorithm evaluates whether the magnetic sublattices can be mapped onto each other through a spatial inversion or lattice translation combined with a spin-flip operation. If such mappings are universally applicable to all magnetic atoms, the material is pure AFM since these symmetry operations guarantee spin degeneracy across the electronic band structure. Conversely, the absence of such inversion or translation symmetries, coupled with the presence of alternative symmetry operations (e.g., rotational or mirror symmetries), indicates altermagnetism. This condition implies the presence of symmetry-protected spin splitting in the band structure despite zero net magnetization. The advantage of amcheck lies in its reliance solely on crystallographic and magnetic structural information, circumventing the need for computationally intensive electronic structure calculations. Consequently, it serves as a rapid, efficient, and theoretically robust tool ideally suited to large-scale, high-throughput screenings. Applying amcheck, we identified 188 potential altermagnetic candidates suitable for further computational examination, while the remaining materials were excluded.
@@
Figure 1: Two-stage high-throughput computational workflow to identify candidate altermagnetic materials (AMs). (a) Initial screening selects experimentally characterized magnetic materials from the MAGNDATA database. Symmetry analysis via amcheck identifies candidate AMs. (b) In-depth computational verification using spin-polarized DFT calculations (VASP, pymatgen, and httk). Materials exhibiting spin splitting ($\geq$50 meV within $\pm$3 eV of $E_{F}$) are confirmed as AM and included in the final database.
@@
In the second screening phase (Fig.1b), input files for spin-polarized self-consistent field (SCF) and non-self-consistent field (non-SCF) density functional theory (DFT) calculations were systematically generated using pymatgen [29] and the High-throughput toolkit (httk) [30]. The workflow manager in httk provides a powerful, flexible automation framework specifically tailored for large-scale computational workflows, significantly enhancing computational efficiency. Importantly, magnetic orders and initial magnetic moments of atoms were directly extracted from MAGNDATA’s experimentally verified .mcif files, ensuring consistency with experimentally established configurations. In cases where the experimentally reported local magnetic moments were particularly small (e.g., below 1 $\mu_{\mathrm{B}}$/atom), we initialized the corresponding magnetic moments to a slightly higher value (typically 1 $\mu_{\mathrm{B}}$) to avoid artificial quenching during self-consistent field (SCF) relaxation.
@@
The resulting spin-polarized SCF calculations, performed using the Vienna Ab-initio Simulation Package [31, 32, 33], were analyzed to identify meaningful spin-splitting. To select only robust altermagnets, a minimum spin-splitting magnitude of 50 meV was used as the selection criterion and this splitting has to occur within an energy window of ±3 eV around the Fermi level. Based on these criteria, 171 materials were validated as exhibiting significant spin-splitting and advanced to the final characterization step, while the remaining candidates were discarded. Detailed electronic band structure calculations were subsequently conducted for these selected 171 altermagnets, enabling direct visualization and confirmation of spin-up and spin-down band splittings. In these computations, the magnetic moments were assumed to be in a collinear configuration, a commonly adopted simplification in high-throughput studies, which is sufficient for identifying the essential altermagnetic characteristics. The present work undertakes an exhaustive and systematic high-throughput screening of the entire MAGNDATA database, unconstrained by specific classes or properties, aiming to identify and comprehensively characterize all potential altermagnetic candidates.
@@
To further quantify the spin splitting, we computed both the average and volumetric spin-splitting metrics. The average spin splitting $\langle\Delta\rangle$ is defined as the mean of the maximum spin splitting across all $\mathbf{k}$-points with eigenstates lying within a 3 eV window around the Fermi level,
@@
Figure 2: Structural, electronic, and spin-resolved properties of UCr2Si2C. (a) Side view of the crystal structure of UCr2Si2C, illustrating the magnetic configuration with arrows representing spin orientations. (b) Corresponding Brillouin zone (BZ) indicating key high-symmetry points. (c) Spin-polarized electronic band structure along high-symmetry paths, where spin-up and spin-down bands are represented by red and blue dashed lines, respectively. (d) Band structures at symmetry-equivalent paths within the BZ, highlighting the inversion of spin-up and spin-down bands due to symmetry operations. (e) Two-dimensional contour plot of spin splitting ($\Delta E$) at $k_{z}=0.5$, showing the plane with maximal spin splitting. (f) The Fermi surface cut at $k_{z}$== 0, through the $\Gamma$-X-M plane, illustrating spin-up (red) and spin-down (blue) contributions. (g–h) Three-dimensional (3D) spin-resolved Fermi surfaces for (g) spin-up and (h) spin-down electronic states, emphasizing spin-dependent characteristics.
@@ III.1 UCr2Si2C
UCr2Si2C crystallizes in a tetragonal structure belonging to the paramagnetic space group $P4/mmm$ (No. 123), characterized by lattice parameters $a=b=3.98\mathrm{\AA }$ and $c=5.16\mathrm{\AA }$ [36]. The structure is composed of alternating uranium and Cr2Si2C layers stacked along the crystallographic $c$-axis. Within each Cr2Si2C layer, chromium atoms form planar Cr2C networks arranged in a checkerboard motif. Magnetic moments are localized exclusively on the Cr atoms, which occupy two symmetry-inequivalent positions: Cr1 at (0.0, 0.5, 0.5)(0.0,\,0.5,\,0.5) and Cr2 at (0.5, 0.0, 0.5)(0.5,\,0.0,\,0.5). These moments are assumed to align antiparallel along the crystallographic $c$ axis, with an estimated magnitude of approximately $0.66\mu_{\mathrm{B}}$.
@@
Under this collinear AFM configuration, UCr2Si2C is associated with the magnetic space group $P4^{\prime}mmm^{\prime}$ (No. 123.343, BNS notation), which determine its magnetic symmetry properties. This group comprises eight symmetry elements: four unitary operations that preserve spin namely, the identity $\{1\,|\,\mathbf{0}\}$, spatial inversion $\{-1\,|\,\mathbf{0}\}$, a twofold rotation about the $z$-axis $\{2_{001}\,|\,\mathbf{0}\}$, and a mirror reflection across the $xy$-plane $\{m_{001}\,|\,\mathbf{0}\}$ as well as four anti-unitary operations combining spatial transformations with time-reversal symmetry. These include the spin-flip twofold rotations $\{2^{\prime}_{100}\,|\,\mathbf{0}\}$ and $\{2^{\prime}_{010}\,|\,\mathbf{0}\}$, and the time-reversed mirror reflections $\{m^{\prime}_{100}\,|\,\mathbf{0}\}$ and $\{m^{\prime}_{010}\,|\,\mathbf{0}\}$. These operations collectively preserve the overall tetragonal crystal symmetry while accommodating a collinear antiferromagnetic configuration with Cr moments aligned along the $c$-axis.
@@
Crucially, although both inversion $\{-1\,|\,\mathbf{0}\}$ and anti-unitary symmetries appear in the group, the combined operation $\mathcal{PT}=\{-1^{\prime}\,|\,\mathbf{0}\}$ does not constitute a symmetry of the system. The absence of $\mathcal{PT}$ symmetry lifts the constraint that would otherwise enforce spin degeneracy, thereby permitting spin splitting even in the absence of spin–orbit coupling. Furthermore, the system lacks the combined spin-rotation and non-primitive lattice translation symmetry $U\tau$, which maps antiparallel magnetic sublattices onto one another via a $180^{\circ}$ spin rotation coupled with a fractional lattice translation. Such symmetries are characteristic of conventional collinear antiferromagnets and protect spin degeneracy throughout the Brillouin zone. In contrast, the symmetry-inequivalent positioning of the Cr atoms in UCr2Si2C precludes the existence of any $U\tau$ operation. The simultaneous breaking of both $\mathcal{PT}$ and $U\tau$ symmetries thus defines the altermagnetic nature of UCr2Si2C.
@@
Figure 2(c) presents the spin-polarized electronic band structure of UCr2Si2C along high-symmetry paths in the Brillouin zone, as defined in Fig.2(b). Spin-up and spin-down states are shown in red and blue, respectively. A prominent feature of the band structure is the coexistence of spin-degenerate and spin-split regions across momentum space, hallmarks of altermagnetic behavior. The maximum energy separation occurs near the $\Delta_{\text{max}}$ point, where the spin splitting reaches approximately 0.70 eV. The average spin splitting across the Brillouin zone is calculated to be 0.31eV, with a volumetric splitting ratio of roughly 47%. To explicitly identify symmetry-driven spin reversals, we analyzed symmetry-equivalent paths in the reciprocal space, as shown in panel (d). Specifically, along paths such as $\Delta_{\mathrm{max}}$$\Gamma$$\Delta^{\prime}_{\mathrm{max}}$ and $ARR^{\prime}$, exchanging the reciprocal-space coordinates $k_{x}\leftrightarrow k_{y}$ directly leads to inversion of the spin polarization (spin-up states become spin-down and vice versa). This momentum-dependent spin reversal arises naturally from the magnetic symmetry operations of the $P4/mmm1^{\prime}$ space group and it distinguishes altermagnetic systems from conventional ferromagnets or Rashba-type spin–orbit coupled systems.
@@
To further visualize the momentum-resolved spin splitting, we plot the two-dimensional (2D) contour map of the spin splitting magnitude ($\Delta E$) on the $k_{z}=0.5$ plane in Fig.2(e). This plane includes the maximum splitting point $\Delta_{\text{max}}=(0.2,0.5,0.5)$, where the maximum spin splitting of approximately 0.70 eV occurs. In this map, red and blue regions respectively indicate positive and negative spin splitting between spin-up and spin-down states. The most intense features appear in the vicinity of $\Delta_{\text{max}}$, with symmetric but sign-reversed spin splitting observed on opposite sides of the Brillouin zone. This antisymmetric pattern of spin splitting, i.e., $\Delta E(\mathbf{k})=-\Delta E(-\mathbf{k})$, is a defining signature of altermagnetic order in this material.

### v2

@@ Abstract
Altermagnetism represents a recently established class of collinear magnetism that combines zero net magnetization with momentum-dependent spin polarization, enabled by symmetry constraints rather than spin-orbit coupling. This distinctive behavior gives rise to sizable spin splitting even in materials composed of light, earth-abundant elements, offering promising prospects for next-generation spintronics applications. Here, we present a comprehensive high-throughput screening of the 2287 entiries comprising the MAGNDATA database, integrating symmetry analysis with spin-polarized density functional theory (DFT) calculations to identify and characterize altermagnetic candidates. Our workflow investigates the collinear structures in the data set and collinear versions of the ones reported to be non-collinear, uncovering 171 materials exhibiting significant spin splitting, spanning both metallic and semiconducting systems. Detailed results are available for all 171 materials, but we also particularly discuss $\mathrm{UCr}_{2}\mathrm{Si}_{2}\mathrm{C}$, $\mathrm{NbMnP}$, and $\mathrm{YRuO}_{3}$ as representative cases with large spin splitting. Furthermore, comparison with the Computational 2D Materials Database (C2DB) and the AiiDA 2D repository gives 9 bulk altermagnets with chemically equivalent 2D counterparts. Crucially, our momentum-resolved analysis reveals that the spin splitting varies strongly across the Brillouin zone, and that the maximal splitting tends to occur away from the high-symmetry paths, a result that directly informs and guides future photoemission experiments. By expanding the catalog of known altermagnets, this work lays a robust foundation for future experimental and theoretical advances in spintronics and quantum materials discovery.
@@ I Introduction
However, discovering new altermagnets with significant spin splitting through trial-and-error experimental approaches is inherently challenging and resource-intensive. To overcome this limitation, high-throughput computational screening, leveraging reliable first-principles density functional theory (DFT) calculations, has emerged as an efficient and powerful methodology. This approach enables the rapid and systematic exploration of large material databases for candidates exhibiting desirable characteristics. Some recent efforts have used the MAGNDATA database for this purpose (see Sec. V for a comparison), yet these studies have largely focused on specific subclasses, such as metallic altermagnets or magnonic excitations in collinear systems, thus leaving much of the altermagnetic landscape unexplored [25, 26]. In this present work, we undertake an exhaustive and systematic high-throughput screening of the entire MAGNDATA [27, 28] database, without constraints to specific classes or properties, aiming to identify and comprehensively characterize all potential altermagnetic candidates. Through this extensive computational investigation, we successfully discover numerous previously unreported altermagnetic materials, validate several known candidates, and provide robust theoretical frameworks for future experimental realizations. This comprehensive approach not only significantly enhances the current understanding of altermagnetic phenomena but also expands the available pool of materials for next-generation spintronic devices, highlighting the critical importance and novelty of this study.
@@ II Screening workflow
We present a rigorous two-stage high-throughput computational screening workflow for the systematic identification of candidate altermagnetic materials, summarized schematically in Fig. 1. In the first stage of screening (Fig. 1a), we retrieved the March 11, 2025 state of the MAGNDATA database hosted by the Bilbao Crystallographic Server [27, 28, 29]. This database contains 2287 published magnetic structures with crystal structures and magnetic configurations primarily obtained via experimental characterization.
@@
These magnetic materials were subsequently analyzed using the recently developed computational tool amcheck [30]. Amcheck operationalizes the symmetry-based theoretical framework underpinning altermagnetism by assessing symmetry operations relating spin-up and spin-down magnetic sublattices. Specifically, the algorithm evaluates whether the magnetic sublattices can be mapped onto each other through a spatial inversion or lattice translation combined with a spin-flip operation. If such mappings are universally applicable to all magnetic atoms, the material is pure AFM since these symmetry operations guarantee spin degeneracy across the electronic band structure. Conversely, the absence of such inversion or translation symmetries, coupled with the presence of alternative symmetry operations (e.g., rotational or mirror symmetries), indicates altermagnetism. This condition implies the presence of symmetry-protected spin splitting in the band structure despite zero net magnetization. The advantage of amcheck lies in its reliance solely on crystallographic and magnetic structural information, circumventing the need for computationally intensive electronic structure calculations. Consequently, it serves as a rapid, efficient, and theoretically robust tool ideally suited to large-scale, high-throughput screenings. The symmetry analysis, both internally in amcheck and for the details we report in the supplementary materials uses Spglib [31, 32].
We used Pymatgen [33] to read the mcif files from MAGNDATA. We applied amcheck to all the collinear structures, which resulted in 157 candidate altermagnetic materials. We then applied amcheck to collinear versions of all noncollinear materials, where we aligned the spins by the algorithm used by Pymatgen when converting a magnetic moment to a scalar: the magnitude of the magnetic moment is retained and the sign assigned by projecting the moment onto a standard axis. The standard axis is arbitrarily chosen to always have the direction defined by the Cartesian vector $(x,y,z)=(1.01,1.02,1.03)$. This resulted in 131 additional candidates. Overall, we obtained a total of 288 potential AM materials from the MAGNDATA database (157 collinear and 131 noncollinear). Eliminating duplicated structures based on chemical composition and space group, the final set was 188 candidates suitable for further computational examination.
@@
Figure 1: Two-stage high-throughput computational workflow to identify candidate altermagnetic materials (AMs). (a) Initial screening selects experimentally characterized magnetic materials from the MAGNDATA database. Symmetry analysis via amcheck followed by deduplication filtering step, yielding a final set of 188 unique AM candidates. (b) In-depth computational verification using spin-polarized DFT calculations (VASP, pymatgen, and httk). Materials exhibiting spin splitting ($\geq$50 meV within $\pm$3 eV of $E_{F}$) are confirmed as AM and included in the final database.
@@
In the second screening phase (Fig. 1b), input files for spin-polarized self-consistent field (SCF) and non-self-consistent field (non-SCF) DFT calculations were systematically generated using Pymatgen and the High-throughput toolkit (httk) [34]. The workflow manager in httk provides a powerful, flexible automation framework specifically tailored for large-scale computational workflows, significantly enhancing computational efficiency. Importantly, magnetic orders and initial magnetic moments of atoms were directly extracted from MAGNDATA’s mcif files, ensuring consistency with experimentally established configurations. In cases where the reported local magnetic moments were particularly small (e.g., below 1 $\mu_{\mathrm{B}}$/atom), we initialized the corresponding magnetic moments to a slightly higher value (typically 1 $\mu_{\mathrm{B}}$) to avoid artificial quenching during SCF relaxation.
@@
The resulting spin-polarized SCF calculations, performed using the Vienna Ab-initio Simulation Package (VASP) [35, 36, 37], were analyzed to identify meaningful spin-splitting. To select only robust altermagnets, a minimum spin-splitting magnitude of 50 meV was used as a selection criterion and this splitting has to occur within an energy window of ±3 eV around the Fermi level. Based on these criteria, 171 materials were validated as exhibiting significant spin-splitting and advanced to the final characterization step, while the remaining candidates were discarded. Detailed electronic band structure calculations were subsequently conducted for these selected 171 altermagnets, enabling direct visualization and confirmation of spin-up and spin-down band splittings. In these computations, the magnetic moments were assumed to be in a collinear configuration, a commonly adopted simplification in high-throughput studies, which is sufficient for identifying the essential altermagnetic characteristics.
@@
To further quantify the spin splitting, we computed both the average and volumetric spin-splitting metrics. The average spin splitting $\langle\Delta\rangle$ is defined as the mean of the maximum spin splitting across all $\mathbf{k}$-points with eigenstates lying within a ±3 eV window around the Fermi level,
@@
Figure 2: Structural, electronic, and spin-resolved properties of UCr2Si2C. (a) Side view of the crystal structure of UCr2Si2C, illustrating the magnetic configuration with arrows representing spin orientations. (b) Corresponding Brillouin zone (BZ) indicating key high-symmetry points. (c) Spin-polarized electronic band structure along high-symmetry paths, where spin-up and spin-down bands are represented by red and blue dashed lines, respectively. (d) Band structures at symmetry-equivalent paths within the BZ, highlighting the inversion of spin-up and spin-down bands due to symmetry operations. (e) Two-dimensional contour plot of spin splitting ($\Delta E$) at $k_{z}=0.5$, showing the plane with maximal spin splitting. (f) The Fermi surface cut at $k_{z}$== 0, through the $\Gamma$-X-M plane, illustrating spin-up (red) and spin-down (blue) contributions. (g–h) Three-dimensional (3D) spin-resolved Fermi surfaces for (g) spin-up and (h) spin-down electronic states, emphasizing spin-dependent characteristics. Note that the absolute orientation of the collinear spin axis does not affect DFT calculations without spin-orbit coupling, hence, in our pictures the spin directions are shown as computed (in +/- z direction) rather than as reported for the respective structures.
@@ III.1 UCr2Si2C
UCr2Si2C crystallizes in a tetragonal structure belonging to the paramagnetic space group $P4/mmm$ (No. 123), characterized by lattice parameters $a=b=3.98\mathrm{\AA }$ and $c=5.16\mathrm{\AA }$ [40]. The structure is composed of alternating uranium and Cr2Si2C layers stacked along the crystallographic $c$-axis. Within each Cr2Si2C layer, chromium atoms form planar Cr2C networks arranged in a checkerboard motif. Magnetic moments are localized exclusively on the Cr atoms, which occupy two symmetry-inequivalent positions: Cr1 at (0.0, 0.5, 0.5)(0.0,\,0.5,\,0.5) and Cr2 at (0.5, 0.0, 0.5)(0.5,\,0.0,\,0.5). These moments are assumed to align antiparallel, with an estimated magnitude of approximately $0.66\mu_{\mathrm{B}}$.
@@
Under this collinear AFM configuration, UCr2Si2C is associated with the magnetic space group $Pm^{\prime}m^{\prime}m$ (No. 47.252, BNS notation), which determine its magnetic symmetry properties. This group comprises eight symmetry elements: four unitary operations that preserve spin namely, the identity $\{1\,|\,\mathbf{0}\}$, spatial inversion $\{-1\,|\,\mathbf{0}\}$, a twofold rotation about the $x$-axis $\{2_{100}\,|\,\mathbf{0}\}$, and a mirror reflection across the plane perpendicular to the $x$-axis $\{m_{100}\,|\,\mathbf{0}\}$ as well as four anti-unitary operations combining spatial transformations with time-reversal symmetry. These include the twofold rotations about the $z$-axis combined with time reversal $\{2^{\prime}_{001}\,|\,\mathbf{0}\}$ and a twofold rotation about the $y$-axis combined with time reversal $\{2^{\prime}_{010}\,|\,\mathbf{0}\}$, and the time-reversed mirror reflections $\{m^{\prime}_{001}\,|\,\mathbf{0}\}$ and $\{m^{\prime}_{010}\,|\,\mathbf{0}\}$. These operations collectively preserve the overall tetragonal crystal symmetry while accommodating a collinear antiferromagnetic configuration with Cr.
@@
Crucially, although both inversion $\mathcal{P}$ $\{-1\,|\,\mathbf{0}\}$ and anti-unitary time revesal $\mathcal{T}=\{-1^{\prime}\,|\,\mathbf{0}\}$ symmetries appear in the group, the combined operation $\mathcal{PT}=\{-1^{\prime}\,|\,\mathbf{0}\}$ does not constitute a symmetry of the system. The absence of $\mathcal{PT}$ symmetry lifts the constraint that would otherwise enforce spin degeneracy, thereby permitting spin splitting even in the absence of spin–orbit coupling. Furthermore, the system lacks the combined spin-rotation and non-primitive lattice translation symmetry $U\tau$, which maps antiparallel magnetic sublattices onto one another via a $180^{\circ}$ spin rotation coupled with a fractional lattice translation. Such symmetries are characteristic of conventional collinear antiferromagnets and protect spin degeneracy throughout the Brillouin zone. In contrast, the symmetry-inequivalent positioning of the Cr atoms in UCr2Si2C precludes the existence of any $U\tau$ operation. The simultaneous breaking of both $\mathcal{PT}$ and $U\tau$ symmetries thus defines the altermagnetic nature of UCr2Si2C.
@@
Figure 2(c) presents the spin-polarized electronic band structure of UCr2Si2C along high-symmetry paths in the Brillouin zone, as defined in Fig.2(b). Spin-up and spin-down states are shown in red and blue, respectively. A prominent feature of the band structure is the coexistence of spin-degenerate and spin-split regions across momentum space, hallmarks of altermagnetic behavior. The maximum energy separation occurs near the $\Delta_{\text{max}}$ point, where the spin splitting reaches approximately 0.72 eV. The average spin splitting across the Brillouin zone is calculated to be 0.31eV, with a volumetric splitting ratio of roughly 47%. To explicitly identify symmetry-driven spin reversals, we analyzed symmetry-equivalent paths in the reciprocal space, as shown in panel (d). Specifically, along paths such as $\Delta_{\mathrm{max}}$$\Gamma$$\Delta^{\prime}_{\mathrm{max}}$ and $ARR^{\prime}$, exchanging the reciprocal-space coordinates $k_{x}\leftrightarrow k_{y}$ directly leads to inversion of the spin polarization (spin-up states become spin-down and vice versa). This momentum-dependent spin reversal arises naturally from the magnetic symmetry operations of the $Pm^{\prime}m^{\prime}m$ space group and it distinguishes altermagnetic systems from conventional ferromagnets or Rashba-type spin–orbit coupled systems.
@@
To further visualize the momentum-resolved spin splitting, we plot the two-dimensional (2D) contour map of the spin splitting magnitude ($\Delta E$) on the $k_{z}=0.5$ plane in Fig. 2(e). This plane includes the maximum splitting point $\Delta_{\text{max}}=(0.2,0.5,0.5)$, where the maximum spin splitting occurs. In this map, red and blue regions respectively indicate positive and negative spin splitting between spin-up and spin-down states. The most intense features appear in the vicinity of $\Delta_{\text{max}}$, with symmetric but sign-reversed spin splitting observed on opposite sides of the Brillouin zone. This antisymmetric pattern of spin splitting, i.e., $\Delta E(\mathbf{k})=-\Delta E(-\mathbf{k})$, is a defining signature of altermagnetic order in this material.